OpenStudy (anonymous):

If \left| ax-1 \right| le 1 where a is a positive even integer, which of the following CANNOT be a value of x^2? A) 0 B) 1/4 C) 1/2 D) 1 E) 4

OpenStudy (anonymous):

sorry the equation is \[\left| ax-1 \right| \le 1\]

OpenStudy (anonymous):

See solve for: \[ax - 1 \le 0\] From here, find x in terms of a.. Are you able to do this?

OpenStudy (anonymous):

sorry there will come 1 in place of 0 in the RHS..

OpenStudy (anonymous):

Why 1?

OpenStudy (anonymous):

Because it is given in the question...

OpenStudy (anonymous):

yea but what do i do after that?

OpenStudy (anonymous):

See the original equation is: \[|ax - 1| \le 1\] See from here we can form two equations: We solve for first: the first equation is: \[ax - 1 \le 1\] Solve this to find x.. try by adding 1 to both the sides..

OpenStudy (anonymous):

after solving you get x ≤ 2/a and x ≤ 0, then what

OpenStudy (anonymous):

No first one is right but second one is incorrect.. Show me how you make the second equation and how you solved it...

OpenStudy (anonymous):

ax-1 ≤ -1 ax ≤ 0 x ≤ 0/a or 0 ?

OpenStudy (anonymous):

Second equation will become: \[-(ax - 1) \le 1\] Now solve for x..

OpenStudy (anonymous):

x ≥ 0/a or 0 then?

OpenStudy (anonymous):

Remember one thing: When we multiply both sides of an Inequality equation by (-1), THEN SIGN OF INEQUALITY REVERSES...

OpenStudy (anonymous):

Yes now you are right.. so what will be the square of x from here??

OpenStudy (anonymous):

greater than 0?

OpenStudy (anonymous):

So what is the value that cannot be possible.. Check from the choices given..

OpenStudy (anonymous):

i would say A but the answer choice says E

OpenStudy (anonymous):

See, find the square of x got in the first equation??

OpenStudy (anonymous):

4/a^2?

OpenStudy (anonymous):

Now combine the x^2: \[0 \le x^2 \le \frac{4}{a^2}\]

OpenStudy (anonymous):

Sorry I have ignored the equal sign there.. So from this: x can take 0 to the values less than 4 because as given "a" is positive even integer: So 4/a^2 will be definitely less than 4.. Agree??

OpenStudy (anonymous):

oooh i get it thank you :D

OpenStudy (anonymous):

Welcome dear..